OpenStudy (anonymous):

What is the surface area of this rectangular prism? A. 99 m2 B. 162 m2 C. 198 m2 D. 81 m2

OpenStudy (mathstudent55):

\(2 \times (LW + LH + WH) \)

OpenStudy (mathstudent55):

\(2 \times (9 ~m \times 6~m + 9 ~m \times 3~ m + 6~m \times 3~m) \)

OpenStudy (anonymous):

198

OpenStudy (mathstudent55):

Correct!!! Great job!